OpenStudy (anonymous):

Wouldn't this be B? Joe earns $40 for playing music on Saturday nights at the local coffee house. He makes an average of $3.25 in tips per hour. Write the function of Joe's earnings, and solve for how much he would make after 5 hours. y = 3.25x + 40; $43.25 y = 3.25x + 40; $56.25 y = 40x + 3.25; $58.25 y = 40x + 3.25; $203.25

OpenStudy (jjwalton04):

well we just fill it in, y=3.25*5+40 3.25*5=16.25 so 16.25+40=56.25 so youre right

OpenStudy (anonymous):

Alright thanks!Just making sure?

OpenStudy (jjwalton04):

ok lol